OpenStudy (candycane):

Hi I'm having a bit of trouble with this question in my math book. My sister and I have worked on it every possiable way we can think of. Here is the question. you are building a rectangular planter for your school garden, You want the area of the bottom to be 90 ft^2. You want the length of the planter to be 3 ft longer then twice its width. What should the dimensions of the bottom of the planter be?

OpenStudy (anonymous):

A=90^2 A=lw w=2(l)

OpenStudy (anonymous):

w=2(l)+3*

OpenStudy (anonymous):

L=15 W=6

OpenStudy (candycane):

Thats the equation my sister and I used, but we couldn't find the answer

OpenStudy (anonymous):

you thought about too hard. do it simple and without all the equation crap

OpenStudy (candycane):

Yea I'm still over thinking it.. I don't get how to solve this at all... Even with out the equation> its still really confusing

OpenStudy (anonymous):

i did that in my head in lik 30sec lol. i did 10*9=90 4*2+3=11 11*4=44.. not enough. 6*2+3=15 15*6=90 PERFECT!

OpenStudy (candycane):

Oh my god that is simple... Thank you so much

OpenStudy (anonymous):

lol, yups. and no problem what so ever. mind giving a medal?

OpenStudy (anonymous):

Let w be the width. Solve the following for w.\[w(2w+3)=90\]